# Ansible Lightspeed: A Code Generation Service for IT Automation

### v1

@@ Abstract.
The availability of Large Language Models (LLMs) which can generate code, has made it possible to create tools that improve developer productivity. Integrated development environments or IDEs which developers use to write software are often used as an interface to interact with LLMs. Although many such tools have been released, almost all of them focus on general-purpose programming languages. Domain-specific languages, such as those crucial for IT automation, have not received much attention. Ansible is one such YAML-based IT automation-specific language. Red Hat Ansible Lightspeed with IBM Watson Code Assistant, further referred to as Ansible Lightspeed, is an LLM-based service designed explicitly for natural language to Ansible code generation.
In this paper, we describe the design and implementation of the Ansible Lightspeed service and analyze feedback from thousands of real users. We examine diverse performance indicators, classified according to both immediate and extended utilization patterns along with user sentiments. The analysis shows that the user acceptance rate of Ansible Lightspeed suggestions is higher than comparable tools that are more general and not specific to a programming language. This remains true even after we use much more stringent criteria for what is considered an accepted model suggestion, discarding suggestions which were heavily edited after being accepted. The relatively high acceptance rate results in higher-than-expected user retention and generally positive user feedback. This paper provides insights on how a comparatively small, dedicated model performs on a domain-specific language and more importantly, how it is received by users.
@@
copyright: none
@@ 1. Introduction
Ansible is a YAML (a plain-text data serialization language) based domain-specific language dedicated to IT automation. A typical Ansible project consists of YAML files, which are organized into playbooks (programs) and roles (libraries). Figure 1 shows an Ansible playbook with it’s sub-components. Playbooks consist of plays, which are a mapping between hosts and the tasks (sequential execution units) that run on those hosts.
Tasks contain a natural language description in the form of a name field, a module name defining the action to execute, and keys (or options) configuring the action.
Figure 1. A typical Ansible playbook structure.
Ansible Lightspeed (RedHat, 2023) is a generative AI service that utilizes the IBM Watson Code Assistant (WCA) to produce code recommendations based on Ansible best practices. WCA for Ansible is an extension of Ansible Wisdom (Pujar et al., 2023), which is a transformer decoder model of 350 million parameters trained on custom natural language, code and Ansible data.
An example of a user interaction with Ansible Lightspeed is presented Figure 2. With Ansible Lightspeed, you can build an Ansible playbook step by step, by providing a natural language description of an Ansible task using its name field. The description becomes a prompt for the model to generate the code of the task.
@@
Code generation models, and some of the systems that utilize them, have emerged as powerful tools for software developers and system analysts (GitHub, 2021; Tab9, 2018; Replit, 2016; Amazon, 2023). Studies show that AI tools for general-purpose programming languages improve productivity and also suggest that novice programmers may benefit more from such tools (Peng et al., 2023; Barke et al., 2023).
Pujar et al. (2023) showed that a relatively small model fine-tuned on high quality Ansible data can outperform a much larger and more general model on Natural Language to Ansible task generation on benchmark data. But to our knowledge, there has been no published study that has involved thousands of users on the use of code assistants for domain-specific languages. Ansible Lightspeed is available since June  2023(RedHat, 2023), and thousands of developers have interacted with the system and consented to provide valuable feedback. In this paper we analyze the interactions and feedback of users of the closed beta and free Tech Preview versions of Ansible Lightspeed, who consented to share their data.
We first analyze usage trends on a temporal basis, to understand when Ansible Lightspeed is used. Next we see if users continue to use Ansible Lightspeed and if so, how regularly. Most importantly, we check to what degree users accept suggestions made by Ansible Lightspeed. This acceptance rate is a crucial comparison metric with other code completion tools in the industry. It is also essential for us to understand how accurate the accepted suggestions were from the user’s perspective. Which is why we perform an edit analysis of the accepted suggestions and come up with a more stringent criteria for acceptance, where we consider a suggestion as accepted only if the user does not edit most or critical parts of the suggestion after accepting it. We analyze and share user feedback of the Ansible Lightspeed service. The key highlights of our analysis are the following:
  • We show that Lightspeed usage is the highest during working hours, on weekdays, suggesting that developers are probably using Ansible Lightspeed for work.
  • Using N-day user retention metric, we show that a Lightspeed user is 4 times more likely than an average Android app user and 3 times more likely than an iOS app user, to keep using Lightspeed after 30 days.
  • We show that the acceptance rate of Ansible Lightspeed is higher than that of similar, but more general, IDE based code completion tools.
  • We develop a much more stringent acceptance criteria and show that even with this stringent criteria, Ansible Lightspeed’s acceptance rate is higher than the simple acceptance rate of other code completion tools.
  • A majority of the users give Ansible Lightspeed 4 and above rating on a scale of 1 to 5, with 5 being the highest rating.
With this analysis we demonstrate how domain expertise, knowledge of language utilization and the application of best practices for system design and user experience can significantly improve recommendation results and user acceptance rates. Given that we have a domain specific model, targeting a well defined use case, we are able to achieve these relatively high user acceptance rates with a relatively small model of size 350 million parameters.
The paper makes the following contributions:
  • A detailed description of the Ansible Lightspeed system and analysis framework (Section 2)
  • An analysis along multiple dimensions of the usage of a service by thousands of users (Section 3).
  • A much more accurate way to measure acceptance rate of model code suggestions by users, by accounting for user edits (Section 3.3).
@@
Figure 2. Ansible Lightspeed’s workflow in the text editor.
@@ 2. Ansible Lightspeed
Figure 3. Ansible Lightspeed application architecture. A: VS Code extension, B: inference pipeline, C: content matching pipeline, D: analysis framework.
@@
  • A content matching pipeline that finds training examples that are similar to the code suggestions.
@@ 2.1. Code Editor Extension
In the pre-processing stage the prompt is checked for valid YAML syntax and the indentations are adjusted based on the requirements of the inference request. After that, it is passed through Ansible Anonymizer (Budder, 2023) to remove any personal, identifiable information. The cleaned up prompt is then passed to the next stage where the inference request is sent to the model server.
@@
IBM Watson Code Assistant for Red Hat Ansible, or WCA-Ansible, is the model that powers Ansible Lightspeed.
WCA-Ansible is a transformer-based decoder trained from scratch on natural language, source code, and Ansible data. We use the Hugging Face Transformers framework (Wolf et al., 2020) to train a 350 million parameters model with a context window of size 1024; this model utilizes fewer parameters than the model available to subscription holders. The model is pre-trained with about 143 billion tokens and then fine-tuned with about 500k natural language to Ansible task samples. The fine-tuning data is created using a technique similar to the one of Ansible Wisdom (Pujar et al., 2023). The pre-training and fine-tuning data is license filtered to remove all data samples coming from files or repositories with restrictive licenses. We also filter the data to remove hate, abuse and profanity. The pre-training data is also deduplicated at a file level. The model server runs on machines containing A10101010 GPUs equipped with sufficient memory to return an inference response within $22$ seconds on average.
@@ 2.3. Content Matching Pipeline
2.4. Analysis Framework
Our data analysis framework uses an event tracking platform to monitor and record a wide range of events, including user interactions, system activities, and user feedback for Ansible Lightspeed. This platform is a critical part of our data collection pipeline, ensuring that we systematically capture important events within our system. It includes user-initiated actions like completion requests, suggestion acceptance, and service feedback, as well as service-related events like server requests, matching sources, and post-processing logs.
The analysis framework consists of two parts: (1) a usage analyser evaluating engagement of the users, and (2) a suggestion analyzer evaluating the quality of the code suggestions.
@@ 2.4. Analysis Framework
3. Analysis
Our analysis aims to provide valuable insights into user behavior of the Ansible Lightspeed service and offer suggestions to improve user engagement and satisfaction. To achieve this, we address the following questions:
  • What are the usage trends of the service?
  • Do users continue to use the service after trying it?
  • Do users accept the suggestions proposed by the service and do they modify them?
  • How do users edit the code after accepting the suggestions?
  • How do users rate their experience using the explicit feedback mechanism?
By answering these questions, we can assess the service’s capacity to engage, satisfy, and retain users, as well as its potential for long-term influence. This multifaceted approach provides a comprehensive understanding of user behavior on the Ansible Lightspeed service and offers valuable insights for enhancing user experience and fostering continued engagement.
Our analysis is based on data collected from users who agreed to share their usage of the closed beta and free Tech Preview versions of the Ansible Lightspeed service from April 15, 2023 to February 4, 2024, a total of 296 days.
3.1. What are the usage trends of the service?
Table 1. Ansible Lightspeed user and usage trend.
Hourly Daily Weekly Monthly
Active user engagements 37 146 490 2.02k
Active users returning over time 35 123 349 796
New users 2 29 163 690
Total predictions 107 1.43k 6.07k 35.2k
Prediction requests from unique users 24 128 425 1.36k
Average prediction requests per user 4.46 11.2 14.3 25.9
Within the domain of evaluating service usability, the usage trend performance indicator plays a crucial role as a quantitative metric (Lewis and Sauro, 2021; Alzahrani et al., 2022). We analyze the total number of active users scrutinized at hourly, daily, weekly, and monthly intervals. Table 1 features varying engagement frequencies and diversification statistics, providing insights into user behavior and engagement patterns.
We are examining the total prediction requests viewed on hourly, daily, and weekly scales, and we notice several interesting patterns. One pattern shows that users tend to make more requests during work hours on weekdays. This increased activity during work hours highlights the service’s usefulness for work-related tasks. Another pattern reveals that some users have a more sporadic usage trend, particularly during weekends and evenings. This suggests that the service is also being used for non-work-related activities, such as personal projects or learning, indicating its versatility and potential to cater to a wide range of user needs and preferences.
The study of ‘Active users returning over time’ (second row of Table 1) provides us with an opportunity for a more in-depth examination of user retention. We expand upon this topic in the next subsection.
3.2. Do users continue to use the service after trying it?
@@
User retention metrics are crucial in app and service analytics, providing valuable insights into user engagement and app longevity (Aspinall et al., 2001). Our study utilizes the N-day User Retention metric to evaluate the Ansible Lightspeed service (Jarvis, 2023). This metric calculates the percentage of users who continue interacting with an app N days after their initial installation or first use, offering a vital assessment of the app’s ability to retain users over time. A high N-day retention percentage indicates that users find lasting value in the app’s features, content, or services, serving as a barometer of user satisfaction, the effectiveness of on-boarding processes, and the overall quality of the app. Figure 5 shows the N-day user retention of Ansible Lightspeed service.
To gain a better understanding of the N-day user retention, ideally we would have liked to compare Ansible Lightspeed with other code completion tools or IDE Plugins. However since N-day user retention is not available for similar tools or for VS Code plugins, we compare it with global data compiled by Business of Apps (Tafradzhiyski, 2023) for Android and iOS apps, as well as their user retention data for 2022 (Handa, 2022) which represents the worldwide average retention rate across 31 mobile app categories. Table 2 presents the results. On Day 1, the average retention rate is 25.3%, gradually decreasing to 5.7% by Day 30. When looking at platform-specific data, iOS has a Day 1 retention rate of 25.65%, while its Day 30 retention rate drops to 4.13%. Android, on the other hand, exhibits a Day 1 retention rate of 23.01%, with its Day 30 retention standing at 2.59%.
These statistics highlight the typical retention patterns in the iOS and Android ecosystems. Retention rates strongly decline over time. However, the Ansible Lightspeed service stands out as an exception to these trends. With a Day 1 retention rate of 44.79%, Lightspeed significantly outperforms global averages and even surpasses the iOS and Android platforms. Additionally, on Day 30, Lightspeed maintains a strong retention rate of 13.66%, demonstrating its ability to retain users over an extended period.
Table 2. Comparison of N-Day user retention of Ansible Lightspeed service with Android and iOS app trends.
Apps 1-Day 3-Day 7-Day 30-Day
Android apps 23.44% 11.30% 7.16% 2.96%
iOS apps 27.06% 13.76% 9.46% 4.40%
Ansible Lightspeed 44.79% 34.17% 33.68% 13.66%
Our findings reveal that Ansible Lightspeed would rank among the top Android applications. Specifically, Ansible Lightspeed falls within the range of the top 100 Android applications regarding user retention data (Chen, 2018). We are the first code completion tool to release N-Day user retention figures and we hope these figures act as baseline for other tools in the future.
$3.$ Do users accept the suggestions proposed by the service and do they modify them?
(a) Distribution of suggestion modifications in the accepted suggestions.
@@
The Ansible Lightspeed service provides inline suggestions in the editor generated from the task name and document content. Users receive these suggestions, which are always multi-line, after entering the task name and moving to the next line. Then, users can either accept the suggestion by pressing the ‘Tab’ key or reject it by pressing the ‘Esc’ key. This interaction is recorded as a telemetry event, and the data is used to evaluate the initial acceptance rate of the model’s suggestions.
We selected a cohort of 3,474 users from a pool of about 6k new users to conduct our user acceptance rate analysis. In the context of our study, we use the term, ”returning user” for any user who used the Ansible Lightspeed service for two or more days. We found 3,474 such returning users, and they have collectively made 74,583 requests to the Ansible Lightspeed service for task suggestions.
Our analysis reveals that out of 74,583 suggestions Lightspeed made, 36,259 suggestions were accepted by users, resulting in an initial acceptance rate of 48.62%.
Table 3. Summary of the user edit analysis.
@@
Total users analyzed 3,474
Total suggestion requests 74,583
Initially accepted suggestions 36,259
Fully accepted suggestions (0% edits) 21,642
Minor edits (¡ 50%) made after acceptance 5,247
Major edits ($\geq$ 50%) made after acceptance 2,641
Deleted suggestions after acceptance 6,433
Module changed in the accepted suggestion 296
To further evaluate these accepted requests, we examined the changes made by users to the accepted suggestions. Figure 6 and Table 3 analyze the edits made to the accepted suggestions.
In 21,642 instances (29.02% of the suggestions and 59.69% of the accepted ones), Ansible Lightspeed’s suggestions perfectly aligned with user requirements. The service accurately predicted the appropriate Ansible module, incorporating the correct options and sub-options, seemingly matching the users’ objectives. Users did not make any modifications to the endorsed suggestion, indicating a high level of satisfaction with the service’s recommendations.
In 5,247 scenarios (7.04% of the suggestions and 14.47% of the accepted ones), users made small adjustments (less than 50% edits) to the accepted suggestions. While Ansible Lightspeed accurately predicted the correct usage of the desired module, users modified one or more module keys (options) and/or values to better align with their specific needs. We examine the edits in greater detail later in the section.
In 2,641 of the remaining suggestions (3.54% of the suggestions and 7.28% of the accepted ones), we observe substantial modifications (50% edits or more) after acceptance, highlighting the crucial need for significant changes in the generated code to meet users’ requirements.
We noticed that users deleted 6,433 accepted suggestions after acceptance, which is 17.74% of the total initially accepted suggestions. Two possible explanations for this behaviour could be:

### v2

# Insights from the Usage of the Ansible Lightspeed Code Completion Service

@@ Abstract.
The availability of Large Language Models (LLMs) which can generate code, has made it possible to create tools that improve developer productivity. Integrated development environments or IDEs which developers use to write software are often used as an interface to interact with LLMs. Although many such tools have been released, almost all of them focus on general-purpose programming languages. Domain-specific languages, such as those crucial for Information Technology (IT) automation, have not received much attention. Ansible is one such YAML-based IT automation-specific language. Ansible Lightspeed is an LLM-based service designed explicitly to generate Ansible YAML given natural language prompt.
This paper first presents the design and implementation of the Ansible Lightspeed service. We then evaluate its utility to developers using diverse indicators, including extended utilization, analysis of user rejected suggestions, as well as analysis of user sentiments. The analysis is based on data collected for 10,696 real users including 3,910 returning users. The code for Ansible Lightspeed service and the analysis framework is made available for others to use.
To our knowledge, our study is the first to involve thousands of users in evaluating code assistants for domain-specific languages. We propose an improved version of user acceptance rate and we are the first code completion tool to present N-Day user retention figures. With our findings we provide insights into the effectiveness of small, dedicated models in a domain-specific context. We hope this work serves as a reference for software engineering and machine learning researchers exploring code completion services for domain-specific languages in particular and programming languages in general.
@@
$c$s: Computing methodologies Natural language processingccs: Software and its engineering Context specific languagesccs: Software and its engineering Application specific development environments
@@ 1. Introduction
Figure 1. A typical Ansible playbook structure consists of playbook definition and tasks list. Ansible Lightspeed generates a task, given the task name and the preceding context. Module name is the first term generated by Lightspeed, followed by the module associated keys and values.
[…]
Ansible (Ansible Community, 2024) is a YAML based domain-specific language dedicated to IT automation. It is one of the most widely used infrastructure as code tools (Guerriero et al., 2019; StackExchange, 2023). Since it is open-source, we do not know the precise number of developers who use Ansible but we can get an indication of its popularity with the Ansible python community package (Team, 2024b) getting over 5 million downloads per month (Flynn, 2024). Ansible also has an active community on GitHub (Team, 2024a) with about 5,000 contributors and about 54,000 commits on a yearly basis, which is likely a small fraction of the actual user base. [^1] Another indication of Ansible’s popularity is from Ansible’s social media accounts of Reddit (Team, 2024c) and X (Team, 2024d), counting 63,000 readers and 72,000 followers, respectively.
@@
A typical Ansible project is organized into playbooks (programs) and roles (libraries). Figure 1 shows an Ansible playbook with it’s sub-components. Playbooks consist of plays, which are a mapping between hosts and the tasks (sequential execution units) that run on those hosts. Tasks contain a natural language description in the form of a name field, a module name defining the action to execute, and keys (or options) configuring the action.
@@
Code generation models, and some of the systems that utilize them, have emerged as powerful tools for software developers and system analysts (GitHub, 2021; Tab9, 2018; Replit, 2016; Amazon, 2023). Studies show that AI tools for general-purpose programming languages improve productivity and also suggest that novice programmers may benefit more from such tools (Weber et al., 2024; Peng et al., 2023; Barke et al., 2023). However, generic models do not perform as well on domain-specific languages. Pujar et al. (2023) showed that a relatively small model fine-tuned on high quality Ansible data can outperform a much larger and more general model on benchmark Natural Language to Ansible-task generation test data. But this work does not involve real users in their development environment interacting with a model.
Red Hat Ansible Lightspeed (RedHat, 2023), further referred to as Ansible Lightspeed or simply Lightspeed, is a generative AI service that utilizes watsonx Code Assistant for Ansible (WCA), an extension of Ansible Wisdom (Pujar et al., 2023), to produce code recommendations based on best practices. An example of a user interaction with Ansible Lightspeed is presented in Figure 2. With Ansible Lightspeed, you can build an Ansible playbook step-by-step, by providing a natural language description of an Ansible task using its name field. The description becomes a prompt for the model to generate the code of the task, which usually consists of the module name followed by module keys and values.
@@
[…]
Figure 2. Ansible Lightspeed’s workflow in the text editor. Users receive Lightspeed suggestions, which are almost always multi-line, after entering the task name and moving to the next line. Then, users can either accept the suggestion by pressing the ‘Tab’ key or reject it by pressing the ‘Esc’ key.
Ansible Lightspeed was available since April 15 2023 to a limited number of users as part of a closed beta release, and then generally to everyone as part of a free Tech Preview version from June 2023  (RedHat, 2023) until May 15 2024. During this period thousands of developers interacted with the system.
In this paper we analyze the interactions and feedback of 10,696 users of the closed beta and free Tech Preview versions of Ansible Lightspeed, who consented to share their data. To our knowledge, there has been no published study, with several thousand users, on the use of code assistants for domain-specific languages.
As a part of our study we analyze usage trends on a temporal basis, to understand when Ansible Lightspeed is used. We see if users continue to use Ansible Lightspeed and if so, how regularly. Most importantly, we check to what degree users accept suggestions made by Ansible Lightspeed. This acceptance rate is a crucial comparison metric with other code completion tools in the industry. It is also essential for us to understand how accurate the accepted suggestions were from the user’s perspective. This is done through an edit analysis of the accepted suggestions. The edit analysis tells us how the user edited the model suggestion before using it. This gives valuable insights on whether the model suggestion was truly useful and clues on how we can improve it. As a result of edit analysis, we also get a more effective and stringent criteria for acceptance, where we consider a suggestion as accepted only if the user does not edit most or critical parts of the suggestion after accepting it. Finally, we analyze and share user qualitative feedback of the Ansible Lightspeed service.
The paper makes the following contributions:
  • A detailed description of the Ansible Lightspeed system and analysis framework. The code for Ansible lightspeed [^2] https://github.com/ansible/ansible-ai-connect-service and the analysis framework [^3] https://github.com/ansible-community/lightspeed-analysis-framework has been made available on GitHub.
  • We propose an improved acceptance criteria that checks if the accepted suggestion was actually used by the user.
  • We show that Ansible Lightspeed’s acceptance rate is higher than that of other, more general, code completion tools.
Given that we have a domain-specific model, targeting a well defined use case, we are able to achieve high user acceptance rates with a relatively small model of size 350 million parameters. Our analysis of user made edits to model suggestions, presents insights on how users use the suggestion and how we can improve it. This study also provides a reference point for software engineering and machine learning researchers investigating the adoption of code completion services, especially for domain-specific languages.
@@ 2. Ansible Lightspeed
[…]
@@
  • A content matching pipeline for attribution, that finds training examples that are similar to the code suggestions and provides the location of the corresponding repositories.
@@ 2.2. Inference Pipeline
In the pre-processing stage the prompt is checked for valid YAML syntax and the indentations are adjusted based on the requirements of the inference request. After that, it is passed through Ansible Anonymizer (Bouder, 2023) to remove any personal, identifiable information. The cleaned up prompt is then passed to the next stage where the inference request is sent to the model server.
@@
IBM watsonx Code Assistant for Red Hat Ansible, or WCA-Ansible, is the model that powers Ansible Lightspeed.
WCA-Ansible is a transformer-based decoder trained from scratch on natural language, source code, and Ansible data. We use the Hugging Face Transformers framework (Wolf et al., 2020) to train a 350 million parameters model with a context window of size 1024; this model utilizes fewer parameters than the model available to subscription holders. The model is pre-trained with about 143 billion tokens and then fine-tuned with about 500k natural language to Ansible task samples. The fine-tuning data is created using a technique similar to the one of Ansible Wisdom (Pujar et al., 2023). The pre-training and fine-tuning data is license filtered to remove all data samples coming from files or repositories with restrictive licenses. We also filter the data to remove hate, abuse and profanity. The pre-training data is also deduplicated at a file level. The model server runs on machines containing A100100100100 GPUs equipped with sufficient memory to return an inference response within $22$ seconds on average.
@@ 2.3. Content Matching Pipeline
2.4. Data Collection and Analysis Framework
Figure 4. Ansible Lightspeed Feedback User Interface. The users can rate their experience using emojis, which we call star-rating. The right most emoji would indicate the best experience or a 5-star rating, and the left most emoji would indicate the worst experience or a 1-star rating. Users are also expected to write a few words to explain their rating in ”Tell us why?” section.
[…]
Ansible Lightspeed collects explicit user feedback, and implicit event tracking. For the explicit feedback, we have created a dedicated user interface (UI) in the sidebar of VS Code Editor as shown in Figure 4. Within this UI, users can provide ratings on a scale of 1 to 5 stars, represented as emojis with rightmost emoji being 5 stars (the best) and leftmost emoji being 1 star (the worst). Along with the rating the users also provide a written feedback on their experience using the service. More on this explicit feedback in Sec. 4.
An event tracking platform monitors and records a wide range of events about user interactions, and system activities. This platform is a critical part of our data collection pipeline, ensuring that we systematically capture important events within our system. It includes user-initiated actions like completion requests, suggestion acceptance, as well as service-related events like server requests, matching sources, and post-processing logs.
The analysis framework consists of two parts: (1) a usage analyzer evaluating engagement of the users, and (2) a suggestion analyzer evaluating the quality of the code suggestions.
@@ 2.4. Data Collection and Analysis Framework
3. Analysis of User Interactions
The goal of our analysis is to determine if the users find the service useful and how we can improve the service to better fit user needs. To achieve this, we try to address the following questions:
  • When do users use Ansible Lightspeed?
  • Do users continue to use the service after trying it once?
  • Do users modify accepted suggestions?
  • At what rate do users accept suggestions?
  • Why do users edit module names in accepted suggestions?
Our analysis is based on data collected from users who agreed to share their usage data of the closed beta and free Tech Preview version of the Ansible Lightspeed service from April 15, 2023, to May 15, 2024, a total of 397 days.
Figure 6. Number of Ansible Lightspeed completion requests per day. Weekdays are displayed in green [ ] and weekends in red [ ]. Average completion requests per day - Monday: 832.69, Tuesday: 1065.08, Wednesday: 990.92, Thursday: 1063.19, Friday: 864.77, Saturday: 241.25, Sunday: 264.90. The data is normalized based on different time zones from where the users made the completion requests.
[…]
3.1. Temporal Analysis
Lightspeed’s goal is to improve the productivity of developers who use Ansible for work. Understanding when requests were made could give some hint that Lightspeed is being used for work related queries.
Figure 6 shows all Lightspeed completion requests [^4] The figures here represent telemetry events, which may contain duplicates. The duplicates will inflate the request count but we believe the trends highlighted below will be remain the same. on a daily basis for the entire analysis period. Requests on weekdays are shown in green and those on weekends are shown in red. One observation is that users tend to make more requests on an average weekday than on an average weekend. We found that ‘Tuesdays’ had the maximum average completion requests per day (average of 1065.08 completion requests) and ‘Saturdays’ had the lowest average completion requests per day (average of 241.25 completion requests). The reason for this weekday and weekend difference could be that a significant chunk of our users are using Lightspeed for work, during weekdays, rather than for study or side projects. However, we need more analysis to be sure.
In Figure 6 we can also see a significant decline in the completion requests between December 23, 2023, and January 12, 2024. This could be due to the Holiday season when many offices are usually closed.
Q: When do users use Ansible Lightspeed? Users make more requests during weekdays, with noticeable drops on weekends and around holidays, which could mean that many users are using Lightspeed for work related requests.
3.2. User Retention
@@
[…]
User retention metrics are crucial in-app and service analytics, providing valuable insights into user engagement and app longevity (Aspinall et al., 2001). Our study utilizes the N-day User Retention metric to evaluate the Ansible Lightspeed service (Jarvis, 2023). This metric calculates the percentage of users who continue interacting with an app N days after their initial installation or first use. Because of this metric’s utility in serving as a barometer for user satisfaction, it is used in many industries, including by iOS and Android app developers (Tafradzhiyski, 2023). For instance, iOS apps have an average Day 30 retention of 4.13%, while the same number for Android apps is 2.59% (Handa, 2022).
Figure 7 shows the N-day user retention of Ansible Lightspeed service. As can be seen, of all the users who use the service on their Day 0, 44.79% return on Day 1, 34.17% on Day 3 and so on. 13.66% users who first used the service 30 days prior, continue to use the service on Day 30.
Based on N-day user retention, we define a returning user, as someone who returned to use the service at least once more after Day 0. Out of a total of about 10,696 unique users, 3,910 or 36.6% were returning users. Since returning users are much more engaged with the service, we assume that they may be using the service for real work, rather than testing it using toy examples. All further analysis is on this cohort of 3,910 are returning users.
Q: Do users continue to use the service after trying it once? On Day 1 the average retention rate is 44.79%. On Day 30 it is 13.66%. 3,910 users out of total 10,696 users, or about 36.5%, are returning users.
3.3. Edit Analysis
(a) Distribution of suggestion modifications in the accepted suggestions. ”Module Changed” indicates Minor edit suggestions in which module was changed.
@@ 3.3. Edit Analysis
[…]
The interactions shown in Figure 2 are recorded as telemetry events, [^5] There are duplicate telemetry events. We can’t be sure why there are duplicate events, but it may be as simple as network issues or user hitting the enter key twice, instead of once. We filter out duplicates and consider only unique requests for our analysis. and this telemetry data is used to calculate the initial acceptance rate of Lightspeed’s suggestions.
We use the data of 3,910 returning users for our user acceptance rate analysis, since we consider returning users to be more engaged. These returning users made a total of 62,099 completion requests to Ansible Lightspeed and received that many suggestions. Out of 62,099 suggestions received by returning users, 40,938 suggestions were accepted, resulting in an initial acceptance rate of 65.92%.
Table 1. Summary of the user edit analysis. A returning user is a user who engaged with the system for more than 1 day. All numbers below, except Total users, are for returning users.
@@
Total users 10,696
Returning users 3,910
Total suggestions 62,099
Initially accepted suggestions 40,938
Fully accepted suggestions (0% edits) 24,811
Minor edits (¡ 50%) made after acceptance 5,672
Major edits ($\geq$ 50%) made after acceptance 2,713
Deleted suggestions after acceptance 7,436
Module changed, but not major edit 306
To understand if these accepted suggestions were used by users, we further examined if changes were made by users to the accepted suggestions. Table 1 and Figure 8.a gives a high level overview of the edits made by returning users on accepted suggestions.
In 24,811 instances (60.60% of the accepted suggestions), Ansible Lightspeed’s suggestions were used without edits so we can assume that they aligned with user requirements. In 5,672 scenarios (13.85% of the accepted suggestions), users made small adjustments, less than 50% edits [^6] Suggestion analyzer part of Sec. 2.4, which explains how edits are calculated., to the accepted suggestions. An example of such a light or minor edit would be, when Ansible Lightspeed accurately predicts the correct module, users modified one or more module keys (options) and/or values to better align with their specific needs.
In 2,713 instances (6.62% of the accepted suggestions), we observe substantial modifications (50% edits or more) after acceptance, indicating that while users accepted the suggestion, they had to heavily modify it before they could actually use it.
We also noticed that users deleted 7,436 suggestions after acceptance, which is 18.16% of the total initially accepted suggestions. Currently there is no way to tell why the users would delete accepted suggestions and move on to next request. One explanation is that users found it challenging to review the suggestion in the greyed text (inline suggestion text) within the VS Code editor, leading them to initially accept and then delete after closer review. Another possible reason could be that they were experimenting with different prompts trying one after another and discarding the answers.
We further analyzed 5,672 instances which were lightly edited, to make sure that they were indeed minor edits. Figure 8.b gives a break down of different kinds of minor edits.
The first observation was that there were 306 instances of minor edits, or 5.4% of all accepted suggestions, in which the module name was also edited. A module could not be considered a minor edit because the keys and values which immediately follow depend on the type of module. If the model gets the module wrong, it is more likely to get rest of the suggestion wrong as well. If the module name is edited, we consider the suggestion to be rejected.
We found that in a majority of remaining instances, the users opted to change only the ‘VALUE’ component of the received suggestions. This involves personalized modifications to variable names, numeric values, and strings, something which maybe hard for Lightspeed to predict based only on the name field prompt and preceding context.
In a much smaller proportion users made changes in both the ‘KEY’ and the ‘VALUE’ parts of the suggested options of the module, added a pair option and value, or removed one. This kind of edit slightly changes the behavior of the task while staying within the context of the suggested module. Instances in which only the ’KEY’ was changed were negligible.
Q: Do users modify accepted suggestions? Yes users modify accepted suggestions. 18.16% of initially accepted suggestions were deleted. 6.62% initially accepted suggestions were heavily modified.
3.4. User Acceptance Rate
The edit analysis described above shows that the acceptance rate as currently defined can be misleading. Which is why we define a new acceptance rate metric, called Edit¡50% acceptance rate. Edit¡50% Acceptance rate is the percentage of accepted suggestions, in which we consider a suggestion accepted only if the following conditions are fulfilled: